Baker Baseball​ Cards, Inc. originally purchased the rookie card of​ Hammerin' Hank Aaron for $ 45.00 After holding the card for 8 ​years, Baker Baseball Cards auctioned the card for $ 450.00

What are the holding period return and the annual return on this​ investment?
Answer #1

Holding Period Return = (Final Value - Initial Value)/Initial value

Holding period return = ($450 - $45)/$45 = 900%

Now, for annual return,

FV = PV * (1 + r)n

450 = 45 * (1 + r)8

10 = (1 + r)8

1.3335 = 1 + r

r = 0.3335

r = 33.35%---> Annual Return